# VeriLLM: A Lightweight Framework for Publicly Verifiable Decentralized Inference

### v2

@@ 5.7 Comparison, Sampling, and On-Chain Adjudication
Comparison and On-Chain Commitment. After all designated verifiers have posted their commitments (or the commitment deadline expires), the Scheduler distributes, off-chain, the inferencer-produced target hidden states $\{S_{i}^{(t)}\}_{t=1}^{T}$ for the relevant segment $i$ to each verifier assigned to that segment. Each verifier $v$ recomputes its local segment to obtain $\{\hat{S}_{i}^{(t)}\}_{t=1}^{T}$ and computes elementwise discrepancies $\Delta^{(t)}=\hat{S}_{i}^{(t)}-S_{i}^{(t)}$. Let $\phi(\cdot)$ denote a scalar aggregation function such as mean absolute error, and $\Psi(\cdot)$ a statistic (mean or clipped mean) summarizing deviations across tokens. The resulting score $T_{v}=\Psi(\{\phi(\Delta^{(t)})\}_{t=1}^{T})$ is compared against a tolerance $\tau$ calibrated for floating-point nondeterminism [51]. The verifier’s binary verdict is
@@
where $r$ is fresh randomness; $C$ is posted during the commit phase, and $(b,r)$ is revealed later [50, 27].
Data Sampling. To ensure verifiers actually perform computation, the contract performs on-chain spot checks of sampled tensor elements [45, 38]. Once verdict commitments are posted, the Scheduler uses a verifiable random function (VRF) to generate unbiased sample indices:
@@
These checks guarantee unbiased sampling, authentic roots, and valid inclusion proofs [44]. Any mismatch can be publicly demonstrated with the Scheduler’s relay signatures, triggering immediate slashing.
@@
where $\varepsilon$ captures tolerated numeric noise [51]. The agreement rate
@@
Tail verifiers ($i=L$) perform an additional token-consistency check [2]:
@@
Calibration experiments provide empirical lower bounds on $\alpha$ and guide the choice of $\varepsilon$ and $\theta$, balancing tolerance and sensitivity [39].
On-Chain Adjudication. After reveal, the contract aggregates revealed verdicts $\{b_{v}\}_{v\in\mathbf{V}}$ from $m$ verifiers. If a majority agree ($B\!\geq\!m/2$), the inference is accepted and honest verifiers share rewards. If most disagree but maintain internal consistency (pairwise hidden-state deviation $\leq\varepsilon$ over at least a $\tau$-fraction of pairs), the negative verdict stands and consistent verifiers are rewarded. Otherwise, the outcome is indeterminate; any verifier may later submit a succinct zero-knowledge proof demonstrating an error, triggering inferencer slashing and bounty payout [40, 38].
@@
This multi-stage design achieves adaptive robustness: typical cases finalize in one verification round with minimal on-chain load, while disputed cases escalate probabilistically until resolved with high confidence [50, 38].
@@ 6.1 Threat Model and Assumptions
VeriLLM operates in a decentralized and adversarial environment where participants may deviate arbitrarily from the protocol to reduce computation cost or bias outcomes. We consider three potentially malicious roles—Inferencers, Verifiers, and a partially trusted or even Byzantine Scheduler. Smart contracts execute deterministically as deployed and are assumed correct by construction [50]. The security goal is to guarantee inference correctness and accountability under these adversarial conditions.
@@
  • EUF-CMA security of digital signatures. Node signatures are existentially unforgeable under chosen-message attack, preventing impersonation or undetected alteration of signed artifacts [45].
  • Collision resistance of hash functions and binding of Merkle commitments. Once a node publishes a Merkle root $\gamma$, no two distinct tensors can produce the same root, ensuring immutability of hidden-state logs [44].
  • Unbiased verifiable randomness. VRF outputs used for role assignment and sampling are unpredictable prior to evaluation and publicly verifiable afterward, precluding adaptive bias [45].
  • Soundness of zero-knowledge proofs. When zero-knowledge proofs attest signature validity or batched inclusion proofs, any false statement is rejected except with negligible probability [38, 39].
@@
  • One-honest-verifier assumption. At least one verifier per task is honest and online. This weaker-than-majority assumption suffices to guarantee safety via full re-execution or sampled audit detection [27].
@@
  • Bounded numeric nondeterminism. Hardware and kernel-induced floating-point variance is modeled as small zero-mean noise within threshold $\varepsilon$; honest executions pass with overwhelming probability [51].
These assumptions align with the threat model of verifiable ML systems such as zkML [38], zkLLM [39], and decentralized inference frameworks [41, 15].
@@ 6.2 Security Objectives
VeriLLM provides cryptographically auditable correctness guarantees for large-scale LLM inference with minimal overhead. When all nodes behave honestly, both the off-chain and on-chain verification phases accept except with negligible probability due to benign numeric noise. If an Inferencer produces an incorrect result—by pruning layers, quantizing weights, truncating decoding, or recycling stale states—then one of three independent mechanisms detects it: (i) full hidden-state recomputation, (ii) commit–then–sample audit enforcing random openings, or (iii) a zero-knowledge proof of inconsistency from an honest verifier [40]. Thus, any materially incorrect inference is provably detectable on chain.
@@
Unpredictable sampling. Sampling indices for audits are derived from VRF outputs computed after all commitments are posted, eliminating adaptivity and ensuring independence between commitments and revealed indices [45].
@@ 6.3 Commit–Then–Sample: Binding and Unpredictability
Together, these properties ensure that a node committing to falsified or incomplete states cannot anticipate future audits and will be detected with overwhelming probability. This commit–then–sample discipline yields deterministic accountability and makes deviation economically irrational in expectation [27, 38].
@@ 6.4.1 Quantization Attack
Thresholds are calibrated from benign full-precision runs. Honest executions pass $T_{\mathrm{off}}$ with probability $1-\operatorname{\textsf{negl}}(\lambda)$, whereas 8-bit quantization causes detectable statistical shifts. (ii) Lightweight on-chain spot check. The contract performs a relaxed test $T_{\mathrm{on}}$ on VRF-sampled final-token scalars, checking $|v_{j}-\mathbf{V}^{(T)}_{j}|\!\leq\!\varepsilon$ and aggregating agreement against a threshold $\theta$ [38]. This sustains liveness while bounding on-chain cost to $O(n)$ sampled entries.
@@ 6.4.2 Early-Termination Attack
Defenses. (i) Tail-token consistency. The last-segment verifier recomputes logits from the reported final hidden state and verifies that the decoding policy yields the claimed token [27]. Any premature termination produces a mismatch at the tail step with probability 1 for deterministic decoding and with high probability for deterministically reseeded stochastic decoding. (ii) Per-step binding. Each segment’s boundary state is individually committed and signed. Modifying terminal states or inserting fake tokens requires forging a signature or producing an inconsistent Merkle opening, both computationally infeasible [44, 45]. Scheduler relay signatures bind token order and prevent state splicing.
@@ 6.4.4 Lazy Verifier (Free-Riding)
Defenses. (i) Commit–then–open discipline. Before seeing sampling indices, each verifier must commit to its final-token tensor root and later open VRF-selected positions with inclusion proofs [45]. Without genuine computation, the probability of producing valid openings is negligible. (ii) Task-type indistinguishability. Because inference and verification share identical RPC interfaces and data formats, workers cannot distinguish verification probes from live inference [27], preventing selective laziness. (iii) Reward gating. Rewards are paid only to verifiers whose openings pass on-chain checks; mismatches or failures trigger stake slashing. Hence, skipping computation is a strictly dominated strategy.
@@ 6.5 Collusion, Censorship, and Dispute
Collusion among verifiers. A coalition may attempt to fabricate mutually consistent hidden states and corresponding Merkle roots such that on-chain openings match internally forged commitments. The commit–then–sample discipline prevents any post hoc adaptation: once a root is posted, the coalition must open VRF-chosen indices consistent with that fixed root. To ensure semantic correctness with respect to the prescribed model, VeriLLM introduces two safeguards: (i) token–state cross-checks at the tail segment, where verifiers decode from recomputed last-layer states and compare tokens to the published outputs; and (ii) a zero-knowledge escalation path, in which any single honest verifier can submit a succinct proof that a challenged prefix of the transcript violates the model’s forward mapping or decoding rule [38, 39, 40]. The contract verifies the proof on-chain, finalizing the inconsistency independently of the coalition’s reports.
@@ 8.1 Main Results
We empirically measured the divergence of hidden states produced during inference across heterogeneous compute devices (tested on Qwen2.5-7B-Instruct [75]; additional models are forthcoming). Although floating-point non-determinism across GPU vendors introduces small numeric drift [76, 77], these deviations are limited in magnitude and statistically distinguishable. Consequently, the hidden-state comparison test can, with high probability, determine whether two traces originate from the same underlying model computation.
We further evaluated replacing full-precision models with quantized variants. Hidden states computed by quantized models exhibit substantially larger—and statistically significant—bias relative to the full-precision baseline, consistent with prior analyses of quantization-induced drift [78, 39]. Hence, simple statistical tests cleanly separate hidden states from normal inference versus those produced under precision degradation or adversarial manipulation.
@@
Bit-aware statistical comparison. Building on these results, we propose a feature-based comparison of hidden states. Given reference $S$ and candidate $\bar{S}$, each floating-point element is decomposed into sign, exponent, and mantissa [77]. With tolerances $e_{w}>0$ (e.g., $e_{w}=10^{-2}$) and $e_{m}\!\in\!\mathbb{N}$ (e.g., 7 mantissa steps), we define:
@@
These rules balance detection sensitivity with tolerance for benign noise and are consistent with prior verifiable ML calibration studies [27, 38].
@@ 8.2 Performance and Overhead Breakdown
We evaluate VeriLLM across five dimensions: throughput and latency, on-chain cost, scalability, ablation, and failure injection [50]. All tests use the 10-node setup described in Section VI.
@@ 8.2.1 Throughput and Latency
For a 7B-parameter model (512-token prompt, 128-token output), total latency per request is 1.37 s: inference 93.1 %, scheduler relay 3.4 %, and verification 3.5 %. Verification overhead averages 0.78 % of total inference time, consistent with lightweight cryptographic pipelines [45]. Increasing verifier parallelism beyond three yields diminishing returns as GPU prefill already saturates compute throughput.
@@ 8.2.2 On-chain Cost Analysis
Each inference produces two transactions: (i) commit (108 k gas, 2.9 kB) and (ii) reveal (134 k gas, 3.2 kB). Merkle proofs require 8–10 hashes per sample, yielding $\approx$ 0.004 USD per verification at current Ethereum L2 prices [79, 80]. VRF verification adds ¡ 4 % of total cost.
@@ 8.2.3 Scalability Experiments
Verification cost per token decreases inversely with verifier count $m$ until m>6m\!>\!6, where on-chain aggregation dominates. Prefill parallelization across $L$ segments achieves near-linear scaling $T_{\text{verify}}\!\approx\!T_{\text{prefill}}/L$, consistent with prior verifiable-inference frameworks [41].
@@ 8.2.5 Failure Injection and Detection Latency
We emulate four attack classes—quantization, early termination, forged output, and lazy verification—via runtime perturbations [27, 39]. Table VI summarizes detection latency (from deviation to slash). All attacks are detected and finalized within 1–3 s, confirming that the commit–sample–reveal protocol yields rapid, tunable defense.

### v3

@@ 5.7 Comparison, Sampling, and On-Chain Adjudication
Comparison and On-Chain Commitment. After all designated verifiers have posted their commitments (or the commitment deadline expires), the Scheduler distributes, off-chain, the inferencer-produced target hidden states $\{S_{i}^{(t)}\}_{t=1}^{T}$ for the relevant segment $i$ to each verifier assigned to that segment. Each verifier $v$ recomputes its local segment to obtain $\{\hat{S}_{i}^{(t)}\}_{t=1}^{T}$ and computes elementwise discrepancies $\Delta^{(t)}=\hat{S}_{i}^{(t)}-S_{i}^{(t)}$. Let $\phi(\cdot)$ denote a scalar aggregation function such as mean absolute error, and $\Psi(\cdot)$ a statistic (mean or clipped mean) summarizing deviations across tokens. The resulting score $T_{v}=\Psi(\{\phi(\Delta^{(t)})\}_{t=1}^{T})$ is compared against a tolerance $\tau$ calibrated for floating-point nondeterminism [rajbhandari2022deep]. The verifier’s binary verdict is
@@
where $r$ is fresh randomness; $C$ is posted during the commit phase, and $(b,r)$ is revealed later [bonneau2015sok, zhang2020delphi].
Data Sampling. To ensure verifiers actually perform computation, the contract performs on-chain spot checks of sampled tensor elements [micali1999verifiable, chen2024zkml]. Once verdict commitments are posted, the Scheduler uses a verifiable random function (VRF) to generate unbiased sample indices:
@@
These checks guarantee unbiased sampling, authentic roots, and valid inclusion proofs [merkle1989certified]. Any mismatch can be publicly demonstrated with the Scheduler’s relay signatures, triggering immediate slashing.
@@
where $\varepsilon$ captures tolerated numeric noise [rajbhandari2022deep]. The agreement rate
@@
Tail verifiers ($i=L$) perform an additional token-consistency check [brown2020language]:
@@
Calibration experiments provide empirical lower bounds on $\alpha$ and guide the choice of $\varepsilon$ and $\theta$, balancing tolerance and sensitivity [sun2024zkllm].
On-Chain Adjudication. After reveal, the contract aggregates revealed verdicts $\{b_{v}\}_{v\in\mathbf{V}}$ from $m$ verifiers. If a majority agree ($B\!\geq\!m/2$), the inference is accepted and honest verifiers share rewards. If most disagree but maintain internal consistency (pairwise hidden-state deviation $\leq\varepsilon$ over at least a $\tau$-fraction of pairs), the negative verdict stands and consistent verifiers are rewarded. Otherwise, the outcome is indeterminate; any verifier may later submit a succinct zero-knowledge proof demonstrating an error, triggering inferencer slashing and bounty payout [qu2025zkgpt, chen2024zkml].
@@
This multi-stage design achieves adaptive robustness: typical cases finalize in one verification round with minimal on-chain load, while disputed cases escalate probabilistically until resolved with high confidence [bonneau2015sok, chen2024zkml].
@@ 6.1 Threat Model and Assumptions
VeriLLM operates in a decentralized and adversarial environment where participants may deviate arbitrarily from the protocol to reduce computation cost or bias outcomes. We consider three potentially malicious roles—Inferencers, Verifiers, and a partially trusted or even Byzantine Scheduler. Smart contracts execute deterministically as deployed and are assumed correct by construction [bonneau2015sok]. The security goal is to guarantee inference correctness and accountability under these adversarial conditions.
@@
  • EUF-CMA security of digital signatures. Node signatures are existentially unforgeable under chosen-message attack, preventing impersonation or undetected alteration of signed artifacts [micali1999verifiable].
  • Collision resistance of hash functions and binding of Merkle commitments. Once a node publishes a Merkle root $\gamma$, no two distinct tensors can produce the same root, ensuring immutability of hidden-state logs [merkle1989certified].
  • Unbiased verifiable randomness. VRF outputs used for role assignment and sampling are unpredictable prior to evaluation and publicly verifiable afterward, precluding adaptive bias [micali1999verifiable].
  • Soundness of zero-knowledge proofs. When zero-knowledge proofs attest signature validity or batched inclusion proofs, any false statement is rejected except with negligible probability [chen2024zkml, sun2024zkllm].
@@
  • One-honest-verifier assumption. At least one verifier per task is honest and online. This weaker-than-majority assumption suffices to guarantee safety via full re-execution or sampled audit detection [zhang2020delphi].
@@
  • Bounded numeric nondeterminism. Hardware and kernel-induced floating-point variance is modeled as small zero-mean noise within threshold $\varepsilon$; honest executions pass with overwhelming probability [rajbhandari2022deep].
These assumptions align with the threat model of verifiable ML systems such as zkML [chen2024zkml], zkLLM [sun2024zkllm], and decentralized inference frameworks [arun2025verde, mei2024fedmoe].
@@ 6.2 Security Objectives
VeriLLM provides cryptographically auditable correctness guarantees for large-scale LLM inference with minimal overhead. When all nodes behave honestly, both the off-chain and on-chain verification phases accept except with negligible probability due to benign numeric noise. If an Inferencer produces an incorrect result—by pruning layers, quantizing weights, truncating decoding, or recycling stale states—then one of three independent mechanisms detects it: (i) full hidden-state recomputation, (ii) commit–then–sample audit enforcing random openings, or (iii) a zero-knowledge proof of inconsistency from an honest verifier [qu2025zkgpt]. Thus, any materially incorrect inference is provably detectable on chain.
@@
Unpredictable sampling. Sampling indices for audits are derived from VRF outputs computed after all commitments are posted, eliminating adaptivity and ensuring independence between commitments and revealed indices [micali1999verifiable].
@@ 6.3 Commit–Then–Sample: Binding and Unpredictability
Together, these properties ensure that a node committing to falsified or incomplete states cannot anticipate future audits and will be detected with overwhelming probability. This commit–then–sample discipline yields deterministic accountability and makes deviation economically irrational in expectation [zhang2020delphi, chen2024zkml].
@@ 6.4.1 Quantization Attack
Thresholds are calibrated from benign full-precision runs. Honest executions pass $T_{\mathrm{off}}$ with probability $1-\operatorname{\textsf{negl}}(\lambda)$, whereas 8-bit quantization causes detectable statistical shifts. (ii) Lightweight on-chain spot check. The contract performs a relaxed test $T_{\mathrm{on}}$ on VRF-sampled final-token scalars, checking $|v_{j}-\mathbf{V}^{(T)}_{j}|\!\leq\!\varepsilon$ and aggregating agreement against a threshold $\theta$ [chen2024zkml]. This sustains liveness while bounding on-chain cost to $O(n)$ sampled entries.
@@ 6.4.2 Early-Termination Attack
Defenses. (i) Tail-token consistency. The last-segment verifier recomputes logits from the reported final hidden state and verifies that the decoding policy yields the claimed token [zhang2020delphi]. Any premature termination produces a mismatch at the tail step with probability 1 for deterministic decoding and with high probability for deterministically reseeded stochastic decoding. (ii) Per-step binding. Each segment’s boundary state is individually committed and signed. Modifying terminal states or inserting fake tokens requires forging a signature or producing an inconsistent Merkle opening, both computationally infeasible [merkle1989certified, micali1999verifiable]. Scheduler relay signatures bind token order and prevent state splicing.
@@ 6.4.4 Lazy Verifier (Free-Riding)
Defenses. (i) Commit–then–open discipline. Before seeing sampling indices, each verifier must commit to its final-token tensor root and later open VRF-selected positions with inclusion proofs [micali1999verifiable]. Without genuine computation, the probability of producing valid openings is negligible. (ii) Task-type indistinguishability. Because inference and verification share identical RPC interfaces and data formats, workers cannot distinguish verification probes from live inference [zhang2020delphi], preventing selective laziness. (iii) Reward gating. Rewards are paid only to verifiers whose openings pass on-chain checks; mismatches or failures trigger stake slashing. Hence, skipping computation is a strictly dominated strategy.
@@ 6.5 Collusion, Censorship, and Dispute
Collusion among verifiers. A coalition may attempt to fabricate mutually consistent hidden states and corresponding Merkle roots such that on-chain openings match internally forged commitments. The commit–then–sample discipline prevents any post hoc adaptation: once a root is posted, the coalition must open VRF-chosen indices consistent with that fixed root. To ensure semantic correctness with respect to the prescribed model, VeriLLM introduces two safeguards: (i) token–state cross-checks at the tail segment, where verifiers decode from recomputed last-layer states and compare tokens to the published outputs; and (ii) a zero-knowledge escalation path, in which any single honest verifier can submit a succinct proof that a challenged prefix of the transcript violates the model’s forward mapping or decoding rule [chen2024zkml, sun2024zkllm, qu2025zkgpt]. The contract verifies the proof on-chain, finalizing the inconsistency independently of the coalition’s reports.
@@ 8.1 Main Results
We empirically measured the divergence of hidden states produced during inference across heterogeneous compute devices (tested on Qwen2.5-7B-Instruct [qwen2_5_technical_report]; additional models are forthcoming). Although floating-point non-determinism across GPU vendors introduces small numeric drift [gupta2015deep, wang2021characterizing], these deviations are limited in magnitude and statistically distinguishable. Consequently, the hidden-state comparison test can, with high probability, determine whether two traces originate from the same underlying model computation.
We further evaluated replacing full-precision models with quantized variants. Hidden states computed by quantized models exhibit substantially larger—and statistically significant—bias relative to the full-precision baseline, consistent with prior analyses of quantization-induced drift [banner2018post, sun2024zkllm]. Hence, simple statistical tests cleanly separate hidden states from normal inference versus those produced under precision degradation or adversarial manipulation.
@@
Bit-aware statistical comparison. Building on these results, we propose a feature-based comparison of hidden states. Given reference $S$ and candidate $\bar{S}$, each floating-point element is decomposed into sign, exponent, and mantissa [wang2021characterizing]. With tolerances $e_{w}>0$ (e.g., $e_{w}=10^{-2}$) and $e_{m}\!\in\!\mathbb{N}$ (e.g., 7 mantissa steps), we define:
@@
These rules balance detection sensitivity with tolerance for benign noise and are consistent with prior verifiable ML calibration studies [zhang2020delphi, chen2024zkml].
@@ 8.2 Performance and Overhead Breakdown
We evaluate VeriLLM across five dimensions: throughput and latency, on-chain cost, scalability, ablation, and failure injection [bonneau2015sok]. All tests use the 10-node setup described in Section VI.
@@ 8.2.1 Throughput and Latency
For a 7B-parameter model (512-token prompt, 128-token output), total latency per request is 1.37 s: inference 93.1 %, scheduler relay 3.4 %, and verification 3.5 %. Verification overhead averages 0.78 % of total inference time, consistent with lightweight cryptographic pipelines [micali1999verifiable]. Increasing verifier parallelism beyond three yields diminishing returns as GPU prefill already saturates compute throughput.
@@ 8.2.2 On-chain Cost Analysis
Each inference produces two transactions: (i) commit (108 k gas, 2.9 kB) and (ii) reveal (134 k gas, 3.2 kB). Merkle proofs require 8–10 hashes per sample, yielding $\approx$ 0.004 USD per verification at current Ethereum L2 prices [chainlinkvrf2023, ethgas2024]. VRF verification adds ¡ 4 % of total cost.
@@ 8.2.3 Scalability Experiments
Verification cost per token decreases inversely with verifier count $m$ until m>6m\!>\!6, where on-chain aggregation dominates. Prefill parallelization across $L$ segments achieves near-linear scaling $T_{\text{verify}}\!\approx\!T_{\text{prefill}}/L$, consistent with prior verifiable-inference frameworks [arun2025verde].
@@ 8.2.5 Failure Injection and Detection Latency
We emulate four attack classes—quantization, early termination, forged output, and lazy verification—via runtime perturbations [zhang2020delphi, sun2024zkllm]. Table VI summarizes detection latency (from deviation to slash). All attacks are detected and finalized within 1–3 s, confirming that the commit–sample–reveal protocol yields rapid, tunable defense.